OpenStudy (anonymous):

A company is manufacturing a gear in the shape of a regular polygon. The measure of each angle of the gear is 162º. How many sides does the gear have?

OpenStudy (anonymous):

\[external \angle=180-162=18\] sum of external angles=360 no. of sides=360/18=20